OpenStudy (anonymous):

Solve algebraically. Clearly define all variables, write an appropriate equation, and solve. One number is 6 more than twice the other number. If the sum of the two numbers is 24, find the two numbers.

OpenStudy (anonymous):

One Number must be 6 more than twice the number and in your case a is your one number

OpenStudy (anonymous):

doesn't fit the above statment

OpenStudy (anonymous):

I thought i was going crazy satellite, or am I? lol

OpenStudy (anonymous):

6=b 18=a

OpenStudy (anonymous):

let's call the the first number n then the other one is 6 more than twice the first one, in other words 2n+6

OpenStudy (anonymous):

^ those answers are correct

OpenStudy (anonymous):

equations would be A=2b+6 A+B=24

OpenStudy (anonymous):

and their sum is 24. so you know that \[n+2n+6=24\]

OpenStudy (anonymous):

making \[3n+6=24\] \[3n=1\] \[n=3\]

OpenStudy (anonymous):

oops, should be \[3n=18\] \[n=3\]

OpenStudy (anonymous):

and you can see that there is clearly more than one way to solve this problem. you can use two variables, or you can use one variable

OpenStudy (anonymous):

n=6 heheh

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

details details...

OpenStudy (anonymous):

also please notice that we could have said one number is n and the other is 24 - n and so \[24-n=2n+6\]

OpenStudy (anonymous):

in other words you have many ways to solve one problem. use two variables, one variable, solve for one in terms of the other in two different ways. the only way you can make a mistake is to divide 18 buy 3 and get 3